neutron stars are some of the densest objects in universe. The form during supernova explosions and very small compared to other astronomical objects you can assume 5*10¹⁷kg/m³ as the average density of a neutron star The earth has a total mass of about 5.97*10²⁴kg what would be the diameter of the earth if it had the density of a neutron star

1
Expert's answer
2022-03-09T15:55:19-0500

"d=\sqrt[3]{\frac{6m}{\pi d}}=0.28~km."